Automobiles arrive at the​ drive-through window at the downtown Baton​ Rouge, Louisiana, post office at the rate of 2 every 10 minutes. The average service time is 1.5 minutes. The Poisson distribution is appropriate for the arrival rate and service times are negative exponentially distributed.

q: If a second​ drive-through window, with its own​ server, were​ added, the average time a car is in the system​ = nothing minutes​ (round your response to two decimal​ places).
Answer #1

Arrival rate, = 2/10 per minute = 0.2 per minute

Average service time = 1.5 minutes

Service rate, = 1/1.5 = 0.667 per minute

/ = 0.2/0.667

= 0.3

q)

Using two drive-through windows (servers) , it is a multiple server (M/M/s) queue model

Refer the above table, and value of / = 0.3, and M = 2

corresponding value of Lq = 0.007

Lq is the average queue length

Average number of cars in system, L = Lq+/ = 0.007+0.3 = 0.307

Average time in system, W = L/

= 0.307/0.2

= 1.535 cars